# Sheared flow profile effects on acoustic impedance eduction in small 3D ducts

### v1

@@ Abstract
We investigate the influence of sheared grazing flow on acoustic propagation in three-dimensional rectangular ducts. We particular focus on small ducts typical of most experimental impedance eduction facilities, for which the flow profile in the duct cross-section varies significantly. We assess the effect of simplifying this inherent two-dimensional flow profile to either a one-dimensional (2D duct) or uniform flow profile. Three flow profiles are considered, namely (i) the tensorised hyperbolic tangent, (ii) the law-of-the-wall, and (iii) a one obtained from a RANS simulation. These flow profiles are used in numerical simulations, based on the solution of the Pridmore-–Brown equation, to perform in silico impedance eduction experiments. Results show that realistic flow profiles can be approximated well in acoustics propagating waves in ducts by uniform or 2D-duct models provided the bulk Mach number is correctly accounted for, which contrasts with previous findings for more simplistic flow profiles. The key conclusion of this work is that if viscous effects are negligible and acoustic impedance is a good representation of a lined wall with grazing flow, then the simplification to a uniform flow is a reasonable approximation and traditional eduction methods are satisfactory accurate.
@@ 1 Introduction
Fan noise is a dominant source in modern turbofan aircraft, and its tonal character makes noise reduction critical for aircraft certification. Acoustic liners, typically a honeycomb core between a rigid backplate and a perforated facesheet, are the main passive treatment. Explicitly modeling the physics of a whole liner remains too expensive due to multiscale interactions — such as acoustic-induced flow and boundary layer — although small-scale high-fidelity simulations have recently become feasible [paduano_impact_2025]. Instead, liners are described by their acoustic impedance, $\tilde{Z}(\omega)=\theta+\mathrm{i}\chi$, with $\theta$ the resistance and $\chi$ the reactance. This frequency-dependent property serves as a boundary condition in acoustic models of aeroengine nacelles.
@@
The impedance of an acoustic liner depends on liner geometry [guess1975Calculation, jones2002effects] and operating conditions such as the Sound Pressure Level (SPL) [murray2012development], grazing flow velocity amplitude [guess1975Calculation], and flow profile [kooi1981Experimental]. Today, impedance eduction is the standard experimental approach for characterizing liners with grazing flow. It infers impedance from in-duct acoustic measurements and a duct propagation model, offering simpler instrumentation, higher repeatability, and broad applicability across liner designs [watson1999validation, jing2008straightforward, elnady2009validation, ferrante2016back, bonomoComparisonSituImpedance2024].
@@
Impedance eduction methods rely on acoustic propagation models for ducts. Common simplifications are that (i) the flow is uniform, and (ii) the Ingard–Myers Boundary Condition (IMBC) [ingard1959influence, myers1980acoustic] can be used to take into account the thin sheared boundary layer effects on acoustic propagation. These simplifications are currently under scrutiny in the scientific community. In principle, the acoustic impedance of liners should remain independent of the incident acoustic field, i.e., one can assume the locally reactive assumption. However, experiments have shown different impedance values when the wave propagation direction changes between upstream- and downstream-propagating waves, suggesting flaws in the modeling used for impedance eduction [renou2011Failure, boden2016Effect, roncen2019influence, boden2017Comparison].
@@
Several attempts have been made to include governing equations, capturing actual liner physics, to solve for this discrepancy. For instance, Ref. [weng2018Flow] found that considering the effects of the viscosity in the acoustic propagation is unable to account for the discrepancy between upstream and downstream and collapse the educed impedances. Ref. [nark2018assessment] found that small changes to the mean flow velocity were able to reduce the gap between the upstream and downstream educed impedances, but not collapse the curves. Ref. [roncen2020WavenumberBased] showed that the difference persists even when considering a three-dimensional inviscid sheared flow profile and a no-slip condition at the lined wall, suggesting that liner physics is not fully captured. Novel boundary conditions have been introduced to better describe liner physics, but none have accurately predicted acoustic wavenumbers when the experimental configuration is changed [spillere2020Experimentally].
@@
Recently, Ref. [bonomoUnpublished] extended the work of nayfeh1974Effect by examining how velocity profile formulations affect impedance eduction. Using the Pridmore–-Brown Equation (PBE) [pridmorebrown1958sound], wavenumbers for realistic profiles were compared against those obtained with the uniform flow assumption and the IMBC. The results showed that the IMBC [ingard1959influence, myers1980acoustic] is adequate for small ducts and low Mach numbers when using realistic profiles, but errors grow when using simplified profiles, increasing with duct width and mean flow velocity. A key limitation of Ref. [bonomoUnpublished] was the restriction to two-dimensional ducts (i.e. 1D flow profiles), leaving the role of three-dimensional effects unresolved. This raises concerns about the conclusions of Ref. [roncen2020WavenumberBased], who employed a hyperbolic tangent profile in their 3D analysis. Since the choice of velocity profile formulation strongly impacts the educed impedance, their findings may be unreliable, and a reassessment with more realistic flow representations is necessary to properly understand sheared flow effects in 3D ducts.
@@
In this paper, we reassess the impact of simplifying the three-dimensional nature of acoustic wave propagation in lined rectangular ducts to either a two-dimensional domain or a uniform flow case, under conditions typical of impedance eduction facilities. Building on our recent findings on the role of axial velocity profile shape [bonomoUnpublished], we revisit earlier conclusions, including those of Ref. [roncen2020WavenumberBased], which relied on simplified flow assumptions. In this paper, we perform virtual in silico experiments using the full 2D version of the PBE as the reference model; using this, we compute the least-attenuated mode wavenumbers for different velocity profiles and then apply the straightforward impedance eduction method to quantify how uniform flow or 2D approximations may distort experimental outcomes.
@@ 2 Governing equations
$(\mathrm{i}\omega+\boldsymbol{u_{0}\cdot\nabla})\left(\dfrac{1}{c_{0}^{2}}(\mathrm{i}\omega+\boldsymbol{u_{0}\cdot\nabla})^{2}\tilde{p}^{\prime}-\nabla^{2}\tilde{p}^{\prime}\right)+2\dfrac{\partial}{\partial z}\big(\boldsymbol{\nabla}\tilde{p}^{\prime}\boldsymbol{\cdot\,\nabla}\!U_{0}\big)=0,$ (1)
@@
$\left(\nabla_{\!\perp}^{2}+\dfrac{\omega^{2}}{c_{0}^{2}}\right)\tilde{p}^{\prime}-k_{z}\left(\dfrac{U_{0}}{\omega}\nabla_{\!\perp}^{2}-\dfrac{2}{\omega}\nabla_{\!\perp}U_{0}\cdot\nabla_{\!\perp}+\dfrac{3\omega U_{0}}{c_{0}^{2}}\right)\tilde{p}^{\prime}-k_{z}^{2}\left(1-\dfrac{3U_{0}^{2}}{c_{0}^{2}}\right)\tilde{p}^{\prime}-k_{z}^{3}\left[\dfrac{U_{0}}{\omega}\left(\dfrac{U_{0}^{2}}{c_{0}^{2}}-1\right)\right]\tilde{p}^{\prime}=0,$ (2)
@@ 2.2 Eigenvalue problem
where the $\mathrm{A}_{j}$ terms involve differentiation in $x$ and $y$ and multiplication by the frequency $\omega$, by the mean flow $U_{0}$ and by its $x$- and $y$-derivatives, and $\boldsymbol{\tilde{p}}^{\prime}$ is the discretized acoustic pressure. In the present work, we follow a pseudo-spectral strategy similar to boyer2011Theoretical, where the problem is discretized by projecting it onto a Gauss–Lobatto grid using Chebyshev polynomials as basis. Finally, to solve the cubic generalized eigenvalue problem given by Eq. (8), auxiliary variables are introduced of the form $\boldsymbol{\tilde{p}}_{p}=k_{z}\boldsymbol{\tilde{p}}_{p-1}$ for $p>0$ [boyd2001Chebyshev, p. 129], and the resulting sparse linear eigenvalue problem is solved using the eigs function from Matlab.
@@
$\dfrac{\mathrm{d}\tilde{p}^{\prime}}{\mathrm{d}x}n_{x}+\dfrac{\mathrm{i}\omega}{Z}\tilde{p}^{\prime}=0,$ (9)
@@
$\dfrac{\mathrm{d}\tilde{p}^{\prime}}{\mathrm{d}x}=0.$ (10)
@@
$\dfrac{\mathrm{d}\tilde{p}^{\prime}}{\mathrm{d}y}=0.$ (11)
In this work, we employed a collocation grid with $N_{x}\times N_{y}=151\times 41$ points in the $x$ and $y$ directions, respectively. The convergence analysis was carried out independently in each direction following the same procedure as Ref. [bonomoUnpublished], yielding an estimated numerical error of $\varepsilon<$1\text{\times}{10}^{-6}$\$. It is worth noting that $N_{y}<N_{x}$ despite $H>W$, owing to the presence of only acoustically hard walls in the $y$ direction.
@@ 3.1 Mean flow profile
The definition of the velocity profile for turbulent flows in a rectangular duct is a non-trivial task. The presence of secondary flows leads to the transport of flow momentum from the center of the duct to the corners, which is then pushed back to the center along the walls to preserve continuity [nakayama1983Calculation]. This creates complex patterns that make obtaining an explicit formulation for the 2D velocity profile challenging. To the best of the authors’ knowledge, no such formulation is currently available in the literature. The common simplification is to adopt a tensorisation of the 1D flow profiles in the $x$ and $y$ directions to obtain a 2D flow profile [roncen2020WavenumberBased]. In this work, we consider three flow profiles for the rectangular duct, two obtained by the traditional tensorisation procedure and one obtained with the assistance of Computational Fluid Dynamics (CFD). The three flow profile considered in this work can be see in Fig. 2.
Figure 2: Flow profiles considered in this work. (a) Tensorized law of the wall; (b) CFD evaluated, and; (c) tensorized hyperbolic tangent. Isolines represent contour axial velocity.
We use the dimensions of the Liner Impedance Test Rig from the Federal University of Santa Catarina (LITR/UFSC), which has a rectangular cross-section with a width of $W=$40\text{\,}\mathrm{mm}$\$ and height $H=$100\text{\,}\mathrm{mm}$\$, as representative of a typical liner testing facility. The frequency range is defined from 500 to $3000\,\mathrm{Hz}$, with steps of $100\,\mathrm{Hz}$. We consider a bulk Mach number of $M=0.3$.
@@
$M(r)=M_{c}\left[\tanh{\left(\dfrac{1-r}{\delta_{t}}\right)}+(1-\tanh{(1/\delta_{t}))\left(\dfrac{1+\tanh{(1/\delta_{t})}}{\delta_{t}}r+(1+r)\right)}(1-r)\right],$ (12)
@@
$\Pi=\dfrac{2}{1+\sqrt{1+4\kappa^{2}{y^{+}_{\text{max}}}^{2}(1-\exp(-y^{+}_{\text{max}}/A^{+})})^{2}}(y^{+}_{\text{max}}-y^{+})\left(\dfrac{y^{+}}{y^{+}_{\text{max}}}\right)^{2},$ (14)
where $y^{+}_{\text{max},x}=Wu_{\tau}/\nu/2$ and $y^{+}_{\text{max},y}=Hu_{\tau}/\nu/2$ are the distances from the wall to the centreline in viscous lengths, for the $x$ and $y$ directions, respectively.
The formulations for the hyperbolic tangent and the turbulent wall law are used to generate the midspan flow profiles in the $x$ and $y$ directions ($y=0$ and $x=0$, respectively). These two 1D flow profiles are tensorised to a 2D flow profile and rescaled for the bulk Mach number by
@@
$M=\int_{-W/2}^{W/2}\int_{-H/2}^{H/2}\dfrac{U_{0}(x,y)}{c_{0}}\mathrm{d}y\mathrm{d}x.$ (16)
@@
$M_{1\mathrm{D}}=\int_{-W/2}^{W/2}\dfrac{U_{0}(x,y=0)}{c_{0}}\mathrm{d}x.$ (17)

### v3

# Realistic sheared flow profile effects on acoustic impedance eduction in small 3D-ducts

@@ Abstract
We investigate the influence of realistic sheared grazing flow on acoustic propagation in three-dimensional rectangular ducts. We show that conclusions reached in the literature about the effects of sheared grazing flow on acoustic propagation in lined ducts are dependent on the flow profiles use in those studies, and that significantly different conclusions are reached once a realistic flow profile is used. We particularly focus on small ducts typical of most experimental impedance eduction facilities, for which velocity gradients are relevant in a significant fraction of the duct cross-section. We assess the effect of simplifying the velocity distribution in the cross-section to either a one-dimensional (2D spanwise-infinite duct) or uniform flow profile. Three flow profiles are considered, namely (i) the tensorised hyperbolic tangent, (ii) the law-of-the-wall, and (iii) a one obtained from a RANS simulation. These flow profiles are used as input in numerical simulations, based on the solution of the 3D-Pridmore-–Brown equation, to perform in silico impedance eduction experiments. Results show that realistic flow profiles can be well approximated for acoustic wave propagation in ducts by uniform or 1D flow profiles, provided the bulk Mach number is correctly accounted for, which contrasts with previous findings considering more simplistic flow profiles. The key conclusion of this work is that if viscous effects are negligible and acoustic impedance is a good representation of a lined wall with grazing flow, then the simplification to a uniform flow is a reasonable approximation and traditional eduction methods are sufficiently accurate.
Nomenclature
$A^{+}$  = Van Driest constant
$c_{0}$  = speed of sound, $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$
$d$  = perforate hole diameter, $\mathrm{m}$
$\boldsymbol{\hat{\mathrm{e}}_{z}}$  = the unitary vector in the $z$ direction
$f$  = frequency, $\mathrm{Hz}$
$f_{c,x}$  = cut-off frequency in the $x$-direction, $\mathrm{Hz}$
$f_{c,y}$  = cut-off frequency in the $y$-direction, $\mathrm{Hz}$
$H$  = duct height, $\mathrm{m}$
$h$  = liner cavity depth, $\mathrm{m}$
$\mathrm{i}$  = imaginary unit, $\sqrt{-1}$
$k_{0}$  = free-field acoustic wavenumber, ${\mathrm{m}}^{-1}$
$k_{s}$  = viscous Stokes wavenumber, ${\mathrm{m}}^{-1}$
$k_{z}$  = axial acoustic wavenumber, ${\mathrm{m}}^{-1}$
$M$  = bulk Mach number
$M_{1\mathrm{D}}$  = midspan (1D) average Mach number
$p^{\prime}$  = acoustic pressure, $\mathrm{Pa}$
$\tilde{p}^{\prime}$  = complex acoustic pressure amplitude, $\mathrm{Pa}$
$r$  = normalized radial coordinate
$R_{\text{cm}}$  = grazing-flow-induced acoustic resistance
$S_{r}$  = nonlinear resistance slope
$S_{m}$  = nonlinear mass reactance
$t$  = perforate facesheet thickness, $\mathrm{m}$
$\boldsymbol{u}_{0}$  = mean axial flow profile, $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$
$U_{0}$  = mean axial flow velocity, $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$
$u_{\tau}$  = friction velocity, $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$
$u_{\tau,x}$  = friction velocity in the $x$-direction, $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$
$u_{\tau,y}$  = friction velocity in the $y$-direction, $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$
$U^{+}$  = normalized mean velocity
$W$  = duct width, $\mathrm{m}$
$x,y,z$  = spatial coordinates, $\mathrm{m}$
$Z$  = normalized acoustic impedance
$Z_{\text{of}}$  = perforate open-face impedance
$\theta$  = normalized acoustic resistance
$\chi$  = normalized acoustic reactance
$\delta$  = boundary-layer thickness, $\mathrm{m}$
$\delta^{*}$  = displacement thickness, $\mathrm{m}$
$\delta_{t}$  = hyperbolic tangent profile shape factor
$\kappa$  = von Kármán constant
$\mu$  = dynamic viscosity, $\mathrm{Pa}\text{\,}\mathrm{s}$
$\nu$  = kinematic viscosity, $\mathrm{m}^{2}\mathrm{/}\mathrm{s}$
$\rho_{0}$  = mean air density, $\mathrm{k}\mathrm{g}\mathrm{/}\mathrm{m}^{3}$
$\sigma$  = perforate open area ratio
$\omega$  = angular frequency, $\mathrm{rad}\text{\,}{\mathrm{s}}^{-1}$
$\xi$  = distance from the wall, $\mathrm{m}$
Subscripts
$c$  = centerline value
$x,y$  = streamwise-normal and transverse directions
$\infty$  = freestream value
Superscripts
++  = normalized viscous units
@@ 1 Introduction
Fan noise is a dominant source in modern turbofan aircraft, and its tonal character makes noise reduction critical for aircraft certification. Acoustic liners, typically a honeycomb core between a rigid backplate and a perforated facesheet, are the main passive treatment. Explicitly modeling the physics of a whole liner remains too expensive due to multiscale interactions — such as acoustic-induced flow and boundary layer — although small-scale high-fidelity simulations have recently become feasible [zhang2016Numerical, paduano_impact_2025]. Instead, liners are described by their acoustic impedance, $\tilde{Z}(\omega)=\theta+\mathrm{i}\chi$, with $\theta$ the resistance and $\chi$ the reactance. This frequency-dependent property serves as a boundary condition in acoustic models of aeroengine nacelles.
@@
The impedance of an acoustic liner depends on liner geometry [guess1975Calculation, jones2002effects] and operating conditions such as the Sound Pressure Level (SPL) [murray2012development], grazing flow velocity amplitude [guess1975Calculation], and flow profile [kooi1981Experimental]. Today, impedance eduction is the standard experimental approach for characterizing liners with grazing flow. It infers impedance from in-duct acoustic measurements and a duct propagation model, offering simpler instrumentation, higher repeatability, and broad applicability across liner designs when compared to the earlier in situ (Dean’s) method. [watson1999validation, jing2008straightforward, elnady2009validation, ferrante2016back, bonomoComparisonSituImpedance2024, rashidi20253D, troian2017Broadband, zhao2025DeepLearningBased].
@@
Impedance eduction methods rely on acoustic propagation models for ducts. Common simplifications are that (i) the flow is uniform, and (ii) the Ingard–Myers Boundary Condition (IMBC) [ingard1959influence, myers1980acoustic] can be used to take into account the thin boundary layer effects on acoustic propagation. These simplifications are currently under scrutiny in the scientific community. In principle, the acoustic impedance of liners should remain independent of the incident acoustic field, under the assumption of local reactivity. However, experiments have shown different impedance values when the wave propagation direction changes between upstream- and downstream-propagating waves, suggesting flaws in the modeling used for impedance eduction [renou2011Failure, boden2016Effect, roncen2019influence, boden2017Comparison]. Traditionally, the upstream/downstream discrepancy has been reported to manifest as a scissor-like behavior of the educed resistance, in which the upstream source tends to yield lower resistance values at low frequencies, and higher resistance values at higher frequencies, than the downstream source, i.e., a crossing of the upstream- and downstream-educed resistance curves as a function of frequency. The crossing point between the two resistance curves is usually observed near the liner resonance
@@
Several attempts have been made to include governing equations, capturing actual liner physics, to solve for this discrepancy. For instance, Ref. [weng2018Flow] found that considering the effects of the viscosity in the acoustic propagation is unable to account for the discrepancy between upstream and downstream and collapse the educed impedances. Ref. [nark2018assessment] found that small changes to the mean flow velocity were able to reduce the gap between the upstream and downstream educed impedances, but not collapse the curves. Ref. [roncen2020WavenumberBased] showed that the difference persists even when considering a three-dimensional inviscid sheared flow profile and a no-slip mean flow boundary condition at the lined wall, suggesting that liner physics is not fully captured. Novel boundary conditions have been introduced to better describe liner physics, but none have accurately predicted acoustic wavenumbers when the experimental configuration is changed [spillere2020Experimentally].
@@
Recently, Ref. [bonomo2026Effect] extended the work of nayfeh1974Effect by examining how velocity profile formulations affect impedance eduction. Using the Pridmore–-Brown Equation (PBE) [pridmorebrown1958sound], wavenumbers for realistic profiles were compared against those obtained with the uniform flow assumption and the IMBC. The results showed that the IMBC [ingard1959influence, myers1980acoustic] is adequate for small ducts and low Mach numbers when using realistic profiles, but errors grow when using simplified profiles, increasing with duct width and mean flow velocity. A key limitation of Ref. [bonomo2026Effect] was the restriction to two-dimensional ducts (i.e. 1D flow profiles), leaving the role of three-dimensional effects unresolved. These observations are consistent with the findings of Ref. [yangShearFlowEffects2024], who used a power-law velocity profile and showed that shear flow has a limited impact on the educed impedance at low Helmholtz numbers (i.e., at low frequencies and in small ducts). Since the choice of velocity profile formulation strongly impacts the educed impedance, the findings of Ref. [roncen2020WavenumberBased] may be unreliable, and a reassessment with more realistic flow representations is necessary to properly understand sheared-flow effects in 3D ducts.
@@
In this paper, we reassess the impact of simplifying the three-dimensional nature of acoustic wave propagation in lined rectangular ducts to either a two-dimensional domain or a uniform flow case, under conditions typical of impedance eduction facilities. Building on our recent findings on the role of axial velocity profile shape [bonomo2026Effect], we revisit earlier conclusions, including those of Ref. [roncen2020WavenumberBased], which relied on simplified flow assumptions. In this paper, we perform virtual in silico experiments using the full 2D version of the PBE as the reference model; using this, we compute the least-attenuated mode wavenumbers for different velocity profiles and then apply the straightforward impedance eduction method to quantify how uniform flow or 2D approximations may distort experimental outcomes. Importantly, the present work demonstrates that the conclusions drawn by Ref. [roncen2020WavenumberBased] regarding the role of sheared flow in three-dimensional ducts warrant reconsideration, as they were obtained using simplified velocity profiles that may not adequately represent a realistic boundary-layer effects. In particular, special attention is given to the role of the effective mean flow Mach number used in the eduction process. As will be shown, ensuring consistency between the cross-section averaged Mach number of the reference configuration and that assumed in the eduction model plays a key role on upstream/downstream discrepancies, and may outweigh the influence of the specific shear profile representation.
@@ 2 Governing equations
$(\mathrm{i}\omega+\boldsymbol{u_{0}\cdot\nabla})\left(\dfrac{1}{c_{0}^{2}}(\mathrm{i}\omega+\boldsymbol{u_{0}\cdot\nabla})^{2}\tilde{p}^{\prime}-\nabla^{2}\tilde{p}^{\prime}\right)\\ +2\dfrac{\partial}{\partial z}\big(\boldsymbol{\nabla}\tilde{p}^{\prime}\boldsymbol{\cdot\,\nabla}\!U_{0}\big)=0,$ (1)
@@
$\left(\nabla_{\!\perp}^{2}+\dfrac{\omega^{2}}{c_{0}^{2}}\right)\tilde{p}^{\prime}-k_{z}\left(\dfrac{U_{0}}{\omega}\nabla_{\!\perp}^{2}-\dfrac{2}{\omega}\nabla_{\!\perp}U_{0}\cdot\nabla_{\!\perp}+\dfrac{3\omega U_{0}}{c_{0}^{2}}\right)\tilde{p}^{\prime}\\ -k_{z}^{2}\left(1-\dfrac{3U_{0}^{2}}{c_{0}^{2}}\right)\tilde{p}^{\prime}-k_{z}^{3}\left[\dfrac{U_{0}}{\omega}\left(\dfrac{U_{0}^{2}}{c_{0}^{2}}-1\right)\right]\tilde{p}^{\prime}=0,$ (2)
@@ 2.1 Uniform flow case
It is emphasized that a uniform mean flow in a straight duct is assumed in the present formulation. Under these conditions, curvature-related terms associated with the Myers boundary condition do not contribute, and no boundary-layer effects are modeled.
@@ 2.2 Eigenvalue problem
where the $\mathrm{A}_{j}$ terms involve differentiation in $x$ and $y$ and multiplication by the frequency $\omega$, by the mean flow $U_{0}$ and by its $x$- and $y$-derivatives, and $\boldsymbol{\tilde{p}}^{\prime}$ is the discretized acoustic pressure. In the present work, we follow a pseudo-spectral strategy similar to boyer2011Theoretical, where the problem is discretized by projecting it onto a Gauss–Lobatto grid using Chebyshev polynomials as basis. Finally, to solve the cubic generalized eigenvalue problem given by Eq. (8), auxiliary variables are introduced to linearize the eigenvalue problem. Specifically, variables of the form $\boldsymbol{\tilde{p}}_{p}=k_{z}\boldsymbol{\tilde{p}}_{p-1}$ for $p>0$ are defined following the standard approach described in [boyd2001Chebyshev, p. 129]. This reformulation yields an equivalent sparse linear eigenvalue problem, which is solved using the eigs function in Matlab.
@@
$\dfrac{\partial\tilde{p}^{\prime}}{\partial x}n_{x}+\dfrac{\mathrm{i}\omega}{c_{0}Z}\tilde{p}^{\prime}=0,$ (9)
@@
$\dfrac{\partial\tilde{p}^{\prime}}{\partial x}=0.$ (10)
@@
$\dfrac{\partial\tilde{p}^{\prime}}{\partial y}=0.$ (11)
In this work, we employed a collocation grid with $N_{x}\times N_{y}=151\times 41$ points in the $x$ and $y$ directions, respectively. The convergence analysis was carried out independently in each direction following the same procedure as Ref. [bonomo2026Effect], yielding an estimated numerical error of $\varepsilon<$1\text{\times}{10}^{-6}$\$. Note that $N_{y}<N_{x}$ despite $H>W$, owing to the presence of only acoustically hard walls in the $y$ direction.
It is also worth noting that early finite-element discretizations of the Pridmore–Brown equation revealed the presence of additional continuous spectra associated with convected vorticity, often referred to as hydrodynamic modes [brambley2009fundamental]. While these modes are not explicitly analyzed in the present work, they form part of the broader theoretical framework of duct acoustics with mean flow.
@@ 3.1 Mean flow profile
The definition of the velocity profile for turbulent flows in a rectangular duct is a non-trivial task. The presence of secondary flows leads to the transport of flow momentum from the center of the duct to the corners, which is then pushed back toward the center along the walls to preserve continuity [nakayama1983Calculation]. This creates complex flow patterns that make obtaining an explicit formulation for the 2D velocity profile challenging. To the best of the authors’ knowledge, no such formulation is currently available in the literature. A common simplification is to adopt a tensorization of 1D flow profiles in the $x$ and $y$ directions to construct a 2D flow profile [roncen2020WavenumberBased]. Here, “tensorized” refers to the construction of a two-dimensional velocity field by combining independent one-dimensional profiles defined along each transverse direction. In this work, we consider three flow profiles for the rectangular duct: two obtained through the traditional tensorization procedure and one obtained with the assistance of Computational Fluid Dynamics (CFD). The three flow profiles considered in this work are shown in Fig. 2, where an experimental mean flow profile measured in the Liner Impedance Test Rig at the Federal University of Santa Catarina (LITR/UFSC) for a bulk Mach number of 0.278 is also included for comparison.
Figure 2: Flow profiles considered in this work: (a) Tensorized law of the wall; (b) CFD evaluated, and; (c) tensorized hyperbolic tangent. (d) Displays the experimental data measured in the LITR/UFSC for a slightly smaller average velocity, for the purposes of comparison. Isolines represent contour axial velocity. (e) shows the comparison of the different profiles for $y=0$ and $x\in[0,W]$, normalized by the maximum velocity.
We use the dimensions of the LITR/UFSC of $W=$40\text{\,}\mathrm{mm}$\$ and height $H=$100\text{\,}\mathrm{mm}$\$, as representative of a typical liner testing facility. The frequency range is defined from 500 to $3000\,\mathrm{Hz}$, with steps of $100\,\mathrm{Hz}$. We consider a bulk Mach number of $M=0.3$. For reference, the cut-off frequencies of the rigid duct are determined by the duct cross-sectional dimensions. In the streamwise-normal ($x$) direction, the first cut-off frequency is given by $f_{c,x}=c_{0}/(2W)$, yielding $f_{c,x}\approx\mbox{$4.29\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$}$. Similarly, in the transverse ($y$) direction, the first cut-off frequency is $f_{c,y}=c_{0}/(2H)$, resulting in $f_{c,y}\approx\mbox{$1.72\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$}$.
@@
[…]
$\left.+\big(1-\tanh{(1/\delta_{t})\big)\!\left(\dfrac{1+\tanh{(1/\delta_{t})}}{\delta_{t}}r+(1+r)\right)}(1-r)\right]\!,$
@@
$\Pi=\dfrac{2(y^{+}_{\text{max}}-y^{+})}{1+\sqrt{1+4\kappa^{2}{y^{+}_{\text{max}}}^{2}(1-\exp(-y^{+}_{\text{max}}/A^{+})})^{2}}\left(\dfrac{y^{+}}{y^{+}_{\text{max}}}\right)^{2}\!,$ (14)
where $y^{+}_{\text{max},x}=Wu_{\tau,x}/\nu/2$ and $y^{+}_{\text{max},y}=Hu_{\tau,y}/\nu/2$ are the distances from the wall to the centreline in viscous lengths, for the $x$ and $y$ directions, respectively. In this work, the friction velocities $u_{\tau,x}$ and $u_{\tau,y}$ were obtained by fitting Eq. (13) to the experimental mean flow data measured at the LITR and shown in Fig. 2d. To estimate $u_{\tau,x}$, experimental data along the line $0\leq x\leq W/2$ at $y=0$ were used, whereas for $u_{\tau,y}$ the fitting was performed using data along $0\leq y\leq H/2$ at $x=0$. This procedure yielded $u_{\tau,x}=$3.95\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$ and $u_{\tau,y}=$3.71\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. This procedure is similar to the one employed in Ref. [bonomo2026Effect] for fitting one-dimensional mean-flow profiles to experimental data.
The formulations for the hyperbolic tangent and the turbulent wall law are used to generate the midspan flow profiles in the $x$ and $y$ directions ($y=0$ and $x=0$, respectively). The shape factor $\delta_{t}$ of the hyperbolic tangent velocity profile was adjusted independently in the streamwise ($x$) and transverse ($y$) directions so that the resulting boundary-layer thickness matches that obtained from the corresponding law-of-the-wall profiles employed in this study. Based on this criterion, the boundary-layer thicknesses of the law-of-the-wall profiles are $\delta_{x}=1.70~\mathrm{mm}$ and $\delta_{y}=3.93~\mathrm{mm}$, which lead to shape factors $\delta_{t}=12.27\%$ and $\delta_{t}=11.34\%$ in the $x$- and $y$-directions, respectively. This choice ensures a consistent comparison between the different mean-flow models and follows the methodology proposed in our recent work Ref. [bonomo2026Effect]. These two 1D flow profiles are tensorised to a 2D flow profile and rescaled for the bulk Mach number by
@@
$M=\dfrac{1}{WH}\int_{-W/2}^{W/2}\int_{-H/2}^{H/2}\dfrac{U_{0}(x,y)}{c_{0}}\mathrm{d}y\mathrm{d}x.$ (16)
@@
$M_{1\mathrm{D}}=\dfrac{1}{W}\int_{-W/2}^{W/2}\dfrac{U_{0}(x,y=0)}{c_{0}}\mathrm{d}x.$ (17)